OpenStudy (anonymous):

If g(x)= 2(x-4), find the value of x if g(x)= 20. A. 32 B. 12 C. 14 D. 10

OpenStudy (anonymous):

just plug them in. you would do like 2(32-4) which is 56 so no.

OpenStudy (anonymous):

So C. 14?

OpenStudy (anonymous):

Is it 14?

jimthompson5910 (jim_thompson5910):

Two ways to do this Method 1: Plug in each answer choice as @WhoaItsJian suggests Method 2: replace g(x) with 20 and solve for x I prefer method 2 g(x)= 2(x-4) 20 = 2(x-4) 20/2 = 2(x-4)/2 10 = x-4 x-4 = 10 x-4+4 = 10+4 x = 14 So x = 14 is a possible answer. We can check by using Method 1 20 = 2(x-4) 20 = 2(14-4) 20 = 2(10) 20 = 20 ... true, so solution is confirmed you are correct @Rebecca_sonnenschein

OpenStudy (anonymous):

Thank you guys(:

OpenStudy (denisaboichuk):

c=14